Calculate the following based on the following information:

Sales: $20 per unit (100,000 units)

Variable costs: $8 per unit

Fixed costs $350,000

a. what is the contribution margin in dollars?

b. what is the contribution margin per unit?

c. what is the contribution margin ratio?
Answer #1

a. Contribution margin = sales - variable cost

= 20*100000 - 8*100000 = $1200000

b. Contribution margin = 20-8 = $12 per unit

c. Contribution margin ratio = contribution margin/ sales

= 12/20 = 60%